OpenStudy (anonymous):

How do I solve this word problem? David has nickles, dimes, and quarters worth $15.75. He has 10 more dimes than nickles and twice as many quarters as dimes. How many of each does he have?

OpenStudy (anonymous):

Think of it this way....q represents how much in quaters...d for dimes and n for nickles. There are three equations here. The first one says the total number of money here is 15.75 so... q + d + n = 15.75 Then ten more dimes than nickles is 10d = n And twice as many quarters than dimes is 2q = d

OpenStudy (anonymous):

So do I have to solve for each of them separately? Thanks, by the way!

OpenStudy (whpalmer4):

I would do it slightly differently: I would let \(q\) be the number of quarters, \(d\) the number of dimes, \(n\) the number of nickels. That means that the value of the coins is given by \[25q+10d+5n\] and as we know the total value is $15.75 (or 1575 cents), our first equation would be \[25q+10d+5n=1575\] Now, we also know that there are 10 more dimes than nickels, so \(d= 10+n\) is another equation. Finally, twice as many quarters as dimes is \(q = 2d\), which is our third equation. We have 3 equations in 3 unknowns, which is sufficient to determine the answer. We can make it into 1 equation in 1 unknown by doing substitutions with the coin relationships: \[25q+10d+5n = 1575\]Use \(q=2d\): \[25(2d)+10d + 5n = 1575\]Now use \(d=10+n\) \[25(2(10+n))+10(10+n)+5n=1575\] Expand that out and solve for \(n\). Once you have that, use the two formulas relating the number of dimes and quarters and nickels to find the values of \(d\) and \(q\).

OpenStudy (anonymous):

Thanks very Much :)

OpenStudy (whpalmer4):

To check your answer, plug d, q and n into 25q + 10d + 5n and make sure the result is 1575.

OpenStudy (kenljw):

25(2(10+n))+10(10+n)+5n=1575 factor (10+n)((25 2+10))+5n=1575 (n+10)12+n=315 divide by 5 13n +120=315 n=15